Pluto is 4 billion miles from the sun. Light from the sun travels at 186000 miles per second. How many hours does it takes for light to reach from the sun to Pluto ?
Answer #1

Given that,

distance b/w sun and pluto, d = 4 * 10^9 miles

speed of light , v = 186000 miles

time t = d / v

t = 4*10^9 / 186000

t = 21505.37 s

t = 21505.37 / 3600 = 5.97 hr

Hence, light takes 5.97 hours  to reach from the sun to Pluto.